Prev Next

SW&IT Meeting Minutes 2020-03-02

Attendees: Carl, David, Igor, Matthieu, Stephen

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200224
LR20200224 - Bec, Matthieu 40.0
LR20200224 - Holmberg, Carl 0.0
LR20200224 - Edmeades, David 40.0
LR20200224 - Sola, Igor 0.0
LR20200224 - Kubanek, Petr 0.0
LR20200224 - Hooper, Stephen 80.0
LR20200224 - Zhang, Yang 40.0
LR20200224 - Zhang, Xianyu 40.0

News from management

  • Director's Review
  • SHARK-VIS PAE/PSR

Previous Action Items

Owner ActionSorted ascending Due Date Completion Date
IS Luci FITS keyword seeing @zenith 2020-02-24  
DE oVirt for PEPSI 2020-02-12  
DE timecapsule obsdev 2020-02-12  

Open topics

  • Return to fifth floor, Mar 5 ?
  • LBTI cancelled / E- time?
  • SSD 2020
  • DX adsec status
  • synology / deeam - status
  • usb3 extender update
  • allsky extender update
  • PK roll-out, remove from g-aliases and IT distribution, keep pkubanek@lbto.org

Team reports

YZ

  • Developing MODS data machine cleaning script.
  • IT#8077 (Appearance of Unique Named MODS files in /newdata and /Repository), in discussion with Olga.
  • Did some setup on lbto@robs machine for 64-bit MODS test run.
  • Together with Olga, we tested MODS programs on the 64-bit robs machine. Some worked (like the GUI), some didn't and require some changing (Python and Perl scripts). I'll work with Olga to change each individual program so that they work on robs. Also compiling a list of things for Stephen to configure/install.
  • Went over the current TCS issues with Petr and Carl.
  • Prepared PCS patch 20200226 (with Carl) for the mountain (not yet activated).
  • Looking into TMS Mathematica program (first trying to get a free Mathematica program).
  • Discussed briefly with Al for AO-Cravat auto-generate data tables (and possibly reports). Next meeting scheduled on March 6.
  • UMAC presentation from Petr on Wednesday

DE

  • Archive backup hiccup, fixed
  • Veeam host license add/version upgrade
    • Still under support for another year+, so we were entitled to the version upgrade with the cheaper purchase
    • Upgrade required significant OS updates; hadn't been updated in 3 years
    • Suspect multiple reboots of the Veeam server got us into memory issues on VMHost1--source of snapshot consolidation errors
      • Seems a host reboot is the way to solve for now
      • Was able to increase stacksize for the operation that was hitting memory limits and causing backups to fail
      • Discovered possible disk failure on VMHost2
        • Monitoring software requires a reboot to function
        • Downtime acceptable today?
    • Upgrade to v.10 successful
    • Ready to set up summit proxy and back up vmhost.mountain
  • Olga decided to postpone laptop OS work; says problem is not occurring as frequently now. Will follow up
  • Projector HDMI audio
    • Can't disable
    • Muted both mic and speaker
  • JH laptop receipt and configure to his specs; waiting for him to use it and comment
  • PK account disables? Is he still contributing?

XZ

  • Last week
    • DX AdSec troubleshoting and tests on the mountain
      • DSP0 board om Crate one was "broken"
      • Crate 1 still not work properly
      • Next step: clean the boards
      • Change the PBP? PBP is in MG.
  • This week
    • Come to the mountain for DX AdSec on Monday
    • Back to Tucson Wednesday night
    • Continue the DX AdSec work

MB

  • Last week
    • OC, WPC, Management
    • UMAC training
    • ROWG #2
    • DMS adsec/allsky wip
    • new SW post is out
  • This week
    • Shark-VIS doco? trip to Italy ?
    • ROWG #2 notes cleanup and send N/A
    • DMS development (websocket, rbj-science ?)
    • OSA tool access to refmem follow up

IS

Last week:
  • PIT
    • Ironed out some issues with the resubmission backend
  • LUCI
    • Preparing Dimm Seeing keyword change
  • OT
    • found a couple of new bugs
    • created a testing document for ISAs

This week:
  • OT
    • Final release?
  • PIT
    • PIT enhancements and release?
    • Start PIT admin backend.
  • LUCI
    • Need to find a working TCS simulator b4 releasing it.

CH

  • Last week:
    • AHS: Refined LBTI AHS client INDI web service polling, added state change logging write to MongoDB.
    • AZCAM: Reset AGW2 azcam for AC
    • TCS: IT review with PK and YZ, patch process review with YZ
    • UMAC: API review with PK

  • This week:
    • AHS: Add write to EPICS for state change to LBTI AHS client. Switch from web service polling to INDI data subscribe?
    • ARGOS: Investigate creating remote GUI client package.

SH

  • Leroy ASC
  • rm507 stability issues
  • IRAF on 64-bit Fedora 30
  • ROWG
  • sensitive data exposure
  • AOeng removed from key list
  • log copy down from mountain

next:

  • Leroy ASC?
  • ovirt?
  • ROWG?
  • telemetry from DMS
  • obs1 slowdown
Topic revision: r12 - 02 Mar 2020, StephenHooper
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback